OpenStudy (mkmkasim):

evaluate ∫1/(x-2)(x^2+1) using partial fractions

OpenStudy (mkmkasim):

i have it set up like this... A/(x-1) +Bx/(x^2+1) +C/(x^2+1)

OpenStudy (anonymous):

\[\frac{A}{x-1}+\frac{Bx+C}{x^2+1}\] is easier

OpenStudy (mkmkasim):

our professor told us to break it up further...

OpenStudy (anonymous):

don't listen you can break it apart later if you need to

OpenStudy (mkmkasim):

alright then lets do it your way lol

OpenStudy (mkmkasim):

wouldnt it be (Bx+C)(x-2) ?

OpenStudy (anonymous):

ok i misread the whole thing it is \[\int \frac{dx}{(x-2)(x^2+1)}\] right?

OpenStudy (mkmkasim):

correct

OpenStudy (anonymous):

ok then \[A(x^2+1)+(Bx+C)(x-2)=1\] that's better

OpenStudy (mkmkasim):

yeah

OpenStudy (anonymous):

now put \(x=2\)

OpenStudy (mkmkasim):

so A would = 5?

OpenStudy (anonymous):

you get \(5A=1\) making \(A=\frac{1}{5}\)

OpenStudy (anonymous):

To check the final result https://www.wolframalpha.com/input/?i=partial+fraction+1%2F%28%28x-2%29%28x^2%2B1%29%29

OpenStudy (anonymous):

you get \(B=-\frac{1}{5}\) right away since you have \(Ax^2+Bx^2=0\) and you know \(A=\frac{1}{5}\) let me know if that is obvious or not

OpenStudy (mkmkasim):

kinda

OpenStudy (anonymous):

mentally multiply this out in your head, don't really multiply it \[A(x^2+1)+(Bx+C)(x-2)=1\] where will you get an \(x^2\) term? one will be from the first term \(Ax^2\) and the other will be the \(Bx^2\) from the second term that is the only place you will get an \(x^2\) so on the left you have \(Ax^2+Bx^2\) and on the right you have no \(x^2\) meaning \[A+B=0\] and since \(A=\frac{1}{5}\) then \(B=-\frac{1}{5}\)

OpenStudy (anonymous):

of course you are free to multiply all that much out, combine like terms, and then equate like coefficients and then solve an annoying system of equations i would look for shortcuts

OpenStudy (anonymous):

similarly, where is the constant going to come from? the \(A\times 1\) in the first term, and the \(-2C\) in the second term so you know \(A-2C=1\) and again since \(A=\frac{1}{5}\) you have \(C=-\frac{2}{5}\)

OpenStudy (mkmkasim):

ohhh. okay gotcha. then you sub those values back into the broken up integral and then solve right?

OpenStudy (anonymous):

yes, check the wolf answer for the nice way to write it i would actually write \[\frac{1}{5}\int \frac{dx}{x-5}-\frac{1}{5}\int\frac{x+2}{x^2+1}dx\]

OpenStudy (anonymous):

first is a log, second is a simple u-sub

OpenStudy (mkmkasim):

okay thanks

OpenStudy (anonymous):

yw